A 500. gram cart moving at .360 m/s, what is its momentum?

If the cart was to bounce off a wall and return with a velocity of -.240 m/s, what is its Impulse be?  




1
Expert's answer
2021-05-05T19:05:50-0400

(a)

"p=mv=0.5\ kg\cdot0.36\ \dfrac{m}{s}=0.18\ kgm/s."

(b)

"J=\Delta p=m(v_f-v_i),""J=0.5\ kg\cdot(-0.24\ \dfrac{m}{s}-0.36\ \dfrac{m}{s})=-0.3\ kgm/s."